# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 2.3. Manifolds, geodesics, and parallel transport
To keep the paper self-contained, we will try to minimize the assumed prerequisite from Lie theory and matrix manifolds.
For the background on matrix manifolds, the reader can consult [5, 7, 1]. In the embedded setup, we consider a smooth subset $\mathcal{M}$ of a matrix space $\mathcal{E}=\mathbb{R}^{n\times m}$, a manifold. We equipped $\mathcal{M}$ with a pseudo-Riemannian metric (also called semi-Riemannian metric) $\mathbf{g}$, which is a smooth scalar-valued bilinear form $\langle,\rangle_{X,\mathbf{g}}$ on the tangent space of $\mathcal{M}$ at each $X\in\mathcal{M}$. We write $\langle,\rangle_{\mathbf{g}}$ if the point $X$ is understood. The bilinear form is assumed to be nondegenerate. When it is positive-definite, we have a Riemannian metric, otherwise the form could have negative but nonzero eigenvalues.
@@
There is more than one way to define covariant derivatives. However, there is a unique connection satisfying two additional conditions, called the Levi-Civita connection, which will be the kind of connection that we consider in this article. The two conditions are torsion-free, or $\Gamma(\mathtt{Y},\mathtt{Z})=\Gamma(\mathtt{Z},\mathtt{Y})$ for two vector fields $\mathtt{Y},\mathtt{Z}$, and metric compatible, that means
(4) $\mathrm{D}_{\mathtt{Y}}\langle\mathtt{Z},\mathtt{Z}\rangle_{\mathbf{g}}=2\langle\mathtt{Z},\nabla_{\mathtt{Y}}\mathtt{Z}\rangle_{\mathbf{g}}.$
A geodesic from $X$ with initial velocity $\xi\in T_{X}\mathcal{M}$ is a curve $\gamma$ on the manifold satisfying the equation $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$, with the initial value condition $\gamma(0)=X,\dot{\gamma}(0)=\xi$. The map $v\mapsto\gamma(1)$ is the Riemannian exponential map, which is not necessarily the same as the Lie group exponential introduced later.
If $\eta\in T_{X}\mathcal{M}$ is another tangent vector at $X$, then the parallel transport of $\eta$ along the geodesic $\gamma(t)$ from 00 to $t$ is the vector $\mathtt{T}_{0}^{t}(\gamma)\eta=\Delta(t)$, where $\Delta$ is a vector field over the curve $\gamma$, that means for each $0\leq s\leq t$, $\Delta(s)$ is tangent to $\mathcal{M}$ at $\gamma(s)$. Additionally, $\Delta$ must satisfy the transport equation
@@
The above discussion also holds for an open subset of a vector space, and the standard differential geometry approach [23] would identify open subsets of an abstract manifold $\mathcal{M}$ with open subsets of a vector space via local coordinates. We can then define a metric on $\mathcal{M}$ by a compatibility requirement on the coordinates. From here, we can define connections and parallel transport without the embedding requirement.
@@
for two vector fields $\mathtt{Y},\mathtt{Z}$, with $\mkern 1.5mu\overline{\mkern-1.5mu\mathtt{Y}\mkern-1.5mu}\mkern 1.5mu$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathtt{Z}\mkern-1.5mu}\mkern 1.5mu$ their horizontal lifts, $\nabla$ and $\mkern 1.5mu\overline{\mkern-1.5mu\nabla\mkern-1.5mu}\mkern 1.5mu$ the Levi-Civita connections on $\mathcal{M}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$, respectively, while $\overline{\left(\nabla_{\mathtt{Y}}\mathtt{Z}\right)}$ is the horizontal lift of $\nabla_{\mathtt{Y}}\mathtt{Z}$, and $\mathrm{H}$ is the projection to the horizontal spaces $\mathcal{H}_{X}$ at each $X\in\mathcal{M}$.
In the next section, we will define matrix groups, a major topic of this paper. It also provides background for quotient manifolds as examples of submersions.
@@ 2.4. Matrix Lie groups, Lie algebra, left-invariant metrics, and matrix quotient spaces
A Lie group is a differentiable manifold with a differentiable group structure. We are interested in subgroups of the generalized linear group $\mathcal{E}=\operatorname{GL}(n)$, which is the (multiplication) group of invertible matrices in $\mathbb{R}^{n\times n}$. Thus, for such a group $\mathrm{G}$, $\operatorname{I}_{n}\in\mathrm{G}$ and $X_{1}^{-1}X_{2}\in\mathrm{G}$ if $X_{1},X_{2}\in\mathrm{G}$.
@@
The tangent space $\mathfrak{g}$ of $\mathrm{G}$ at $\operatorname{I}=\operatorname{I}_{n}$ is a Lie subalgebra of $\mathbb{R}^{n\times n}$. Recall a vector subspace $\mathfrak{g}$ of the square matrice space $\mathcal{E}=\mathbb{R}^{n\times n}$ is called a Lie subalgebra if it is closed under the Lie bracket, that means for two matrices $\mathtt{a},\mathtt{b}$ in $\mathfrak{g}$,
(7) $[\mathtt{a},\mathtt{b}]:=\mathtt{a}\mathtt{b}-\mathtt{b}\mathtt{a}\subset\mathfrak{g}.$
@@
The Lie bracket is anticommutative and satisfies Jacobis identity [7, Chapter 3]. The matrix exponential map $\operatorname{expm}$ maps $\mathfrak{g}$ to $\mathrm{G}$. Conversely, given a Lie subalgebra $\mathfrak{g}$ of $\mathbb{R}^{n\times n}$, we can construct a Lie group with Lie algebra $\mathfrak{g}$. In this instant, $\operatorname{expm}$ corresponds to the Lie exponential map of $\operatorname{GL}(n)$, and its subgroups. We have
(8) $\operatorname{expm}(A)X\operatorname{expm}(-A)=\operatorname{expv}(\operatorname{ad}_{A},X),$
@@
where $\operatorname{ad}_{A}:X\mapsto[A,X]$ and $\operatorname{expv}$ is the exponential action in section 2.1 [7, Chapter 3]. From this, $\operatorname{expm}(A)X\operatorname{expm}(-A)\in\mathfrak{g}$ if $A,X\in\mathfrak{g}$. The tangent space of $\mathrm{G}$ at $X\in\mathrm{G}$ could be identified with $X\mathfrak{g}$, or $\xi\in T_{X}\mathrm{G}$ if and only if $X^{-1}\xi\in\mathfrak{g}$. From here, if $\langle$,$\rangle$ , ⟩ is a nondegenerate form on $\mathfrak{g}$, we can define a pseudo-Riemannian metric on $\mathrm{G}$ by defining for two tangent vectors $\xi,\eta$ at $X$
(9) $\langle\xi,\eta\rangle_{X}=\langle X^{-1}\xi,X^{-1}\eta\rangle,$
which gives $\mathrm{G}$ a left-invariant pseudo-Riemannian metric, as $\langle X_{1}\xi,X_{1}\eta\rangle_{X_{1}X}=\langle\xi,\eta\rangle_{X}$ for $X_{1}\in\mathrm{G}$.
In applications here, we will consider right actions, smooth maps ¯×$\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\times\mathrm{K}\mapsto\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$, maping $(X,K)\in\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\times\mathrm{K}$ to $XK$ that are compatible with group multiplication from the right. The actions are assumed proper such that the relationship $X\sim XK$ defines a manifold structure on the quotient $\mathcal{M}/\mathrm{K}$. This is satisfied for the case of compact groups, which we consider here. In this case, we can use the notion of horizontal lift above to compute connections, geodesics, and parallel transports.
2.5. Notations
The generalized linear group $\operatorname{GL}(n)$ is the group of all invertible matrices in $\mathbb{R}^{n\times n}$, the connected subgroup with positive determinant is $\operatorname{GL}^{+}(n)$, the corresponding Lie algebra $\mathfrak{gl}(n)$ could be identified with $\mathbb{R}^{n\times n}$. The special orthogonal group $\mathrm{SO}(n)$ is the group of orthogonal matrices of determinant $1$, its Lie algebra $\mathfrak{so}(n)$ could be identified as a vector space with the space of antisymmetric matrices $\mathrm{Skew}_{n}$. Its orthogonal complement is the space $\mathrm{Sym}_{n}$ of symmetric matrices. The Stiefel, flag and Grassmann manifolds with notations $\mathrm{St}_{n,p},\mathrm{Gr}_{n,p},\operatorname{Flag}(\vec{d})$ will be introduced later. We write $\mathrm{sym}$ for the symmetrize operator $A_{\mathrm{sym}}=\frac{1}{2}(A+A^{\mathsf{T}})$ for $A\in\mathbb{R}^{n\times}$ and $\mathrm{skew}$ for the antisymmetrize operator $A_{\mathrm{skew}}=\frac{1}{2}(A-A^{\mathsf{T}})$.
3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
We now turn to bilinear forms on Lie subalgebras and left-invariant metrics on the corresponding groups.
@@ 3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
The trace form is the bilinear map $(\mathtt{a},\mathtt{b})\mapsto\operatorname{Tr}\mathtt{a}\mathtt{b}=:\langle\mathtt{a},\mathtt{b}\rangle_{\mathcal{K}}$ (Green the subscript $\mathcal{K}$ is in honor of Killing, who defined the closely related Killing form). Using the commutative property of trace, we have
(10) $\operatorname{Tr}[\mathtt{a},\mathtt{b}]\mathtt{c}=\operatorname{Tr}\mathtt{b}[\mathtt{c},\mathtt{a}],$
@@
or the trace form is the cyclic invariant. As any strictly upper-triangular matrix $\mathtt{a}$ has $\operatorname{Tr}\mathtt{a}^{2}=0$, the trace form is not positive definite. A symmetric matrix $\mathtt{a}$ has $\operatorname{Tr}\mathtt{a}^{2}\geq 0$, while an antisymmetric matrix has $\operatorname{Tr}\mathtt{a}^{2}\leq 0$. However, the trace form has a very nice property that if it is nondegenerate on $\mathfrak{g}$, it induces a bi-invariant pseudo-Riemannian metric on the corresponding group with closed-form geodesic.
@@
The Frobenius form, or Frobenius inner product $(\mathtt{a},\mathtt{b})\mapsto\langle\mathtt{a},\mathtt{b}\rangle_{\mathcal{F}}=:\operatorname{Tr}\mathtt{a}\mathtt{b}^{\mathsf{T}}$ is positive-definite. However, it may not induce a metric with closed-form geodesics. We will focus on subspaces that transpose to themselves, where we can use the nice properties of trace and Frobenius forms to construct tractable metrics.
@@
Transposable Lie subalgebras have an important property. Let us explain it through an example of a space of block matrices. For $k<n$, let $\mathfrak{a}$ be a copy of $\mathbb{R}^{k\times k}$, considered as a subspace of $\mathcal{E}=\mathbb{R}^{n\times n}$ by identifying $\mathfrak{a}$ with the space of upper-left $k\times k$ block diagonal matrices, and $\mathfrak{a}_{{}_{\perp}}$ its Frobenius orthogonal complements. Thus $\mathfrak{a}$ consists of matrice of the form $\begin{bmatrix}A&0\\ 0&0\end{bmatrix}$ with $A\in\mathbb{R}^{k\times k}$, while $\mathfrak{a}_{{}_{\perp}}$ consists of those of the form $\begin{bmatrix}0&B\\ C&D\end{bmatrix}$, with $D\in\mathbb{R}^{(n-k)\times(n-k)},B\in\mathbb{R}^{k\times(n-k)},C\in\mathbb{R}^{(n-k)\times k}$. Notice that $\mathfrak{a}$ and $\mathfrak{d}=\{\begin{bmatrix}0&0\\ 0&D\end{bmatrix}|D\in\mathbb{R}^{(n-k)\times(n-k)}\}$ commutes, while the Lie bracket of $\mathfrak{a}$ or $\mathfrak{d}$ with a matrix of the form $\begin{bmatrix}0&B\\ C&0\end{bmatrix}$ is another matrix of that form. All three spaces $\mathfrak{a},\mathfrak{d}$ and the remaining part are transposable. The splitting of $\mathbb{R}^{n\times n}$ to these three parts is a special case of proposition 1 below where $\mathfrak{g}$ is $\mathbb{R}^{n\times n}$, $\mathfrak{a}$ is the space of top $k\times k$ block matrices and $\mathfrak{d}$ is $\mathfrak{a}_{{}_{\top}}$, the space of bottom $(n-k)\times(n-k)$ matrices. The remaining space with zero diagonal blocks is $\mathfrak{a}_{{}_{\Join}}$. We can add $\mathfrak{a}_{{}_{\Join}}$ to either $\mathfrak{a}$ or $\mathfrak{a}_{{}_{\top}}$ to form the orthogonal complement of the remaining algebra.
@@
Assume $\mathfrak{g}$ is transposable. Then the trace form is nondegenerate on $\mathfrak{g}$. Let $\mathfrak{v}\subset\mathfrak{g}$ be a transposable subspace. Let $\mathfrak{v}_{{}_{\perp}}$ be the orthogonal complement of $\mathfrak{v}$ under the trace form $\mathcal{K}$
(12) $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{h}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{h}\mathfrak{v}=\{0\}\}.$
Then $\mathfrak{v}_{{}_{\perp}}$ is also transposable and is the orthogonal complement under the Frobenius form $\mathcal{F}$, and vice-versa. The projection $\mathtt{p}_{\mathfrak{v}}$ of $\mathfrak{g}$ onto $\mathfrak{v}$ under $\mathcal{K}$ and $\mathcal{F}$ is identical and commutes with transpose, $\mathtt{p}_{\mathfrak{v}}(\mathtt{h}^{\mathsf{T}})=\mathtt{p}_{\mathfrak{v}}(\mathtt{h})^{\mathsf{T}}$ for $\mathtt{h}\in\mathfrak{g}$.
If $\mathfrak{a}$ is a transposable subalgebra, then we have
[…]
and $\mathfrak{a}_{{}_{\Join}}$ is also transposable. The space $\mathfrak{a}_{{}_{\top}}$ defined by either one of the equations below is a transposable Lie subalgebra of $\mathfrak{g}$
(14) $\mathfrak{a}_{{}_{\top}}=\{\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}|\operatorname{Tr}\mathtt{c}[\mathfrak{a},\mathfrak{a}_{{}_{\perp}}]=\{0\}\},$
(15) $\mathfrak{a}_{{}_{\top}}=\{\mathtt{c}\in\mathfrak{a}_{{}_{\perp}}|[\mathtt{c},\mathfrak{a}]=\{0\}.$
The correspondence $\mathfrak{a}\mapsto\mathfrak{a}_{{}_{\top}}$ is involutive, $(\mathfrak{a}_{{}_{\top}})_{{}_{\top}}=\mathfrak{a}$. We have $[\mathfrak{a},\mathfrak{a}_{{}_{\Join}}]\subset\mathfrak{a}_{{}_{\Join}}$ and $[\mathfrak{a}_{{}_{\top}},\mathfrak{a}_{{}_{\Join}}]\subset\mathfrak{a}_{{}_{\Join}}$ and we have an orthogonal decomposition under either $\mathcal{K}$ or $\mathcal{F}$
(16) $\mathfrak{g}=\mathfrak{a}\oplus\mathfrak{a}_{{}_{\Join}}\oplus\mathfrak{a}_{{}_{\top}}.$

### v2

@@ 2.2. Manifolds, geodesics, and parallel transport
For the background on matrix manifolds, the reader can consult [7, 9, 1]. We will also need background on naturally reductive homogeneous spaces, see [33], reviewed in this section. In the embedded setup, we consider a smooth subset $\mathcal{M}$ of a matrix space $\mathcal{E}=\mathbb{R}^{n\times m}$, a manifold. We equip $\mathcal{M}$ with a pseudo-Riemannian metric (also called semi-Riemannian metric) $\mathbf{g}$, which is a smooth scalar-valued bilinear form $\mathbf{g}_{X}\langle,\rangle$ on the tangent space of $\mathcal{M}$ at each $X\in\mathcal{M}$. We write $\mathbf{g}\langle,\rangle$ or $\langle,\rangle_{\mathbf{g}}$ if the point $X$ is understood. The bilinear form is assumed to be nondegenerate. When it is positive-definite, we have a Riemannian metric, otherwise, the form could have negative but nonzero eigenvalues.
@@
There is more than one way to define covariant derivatives. However, there is a unique connection satisfying two additional conditions, called the Levi-Civita connection, which will be the kind of connection that we consider in this article. The two conditions are torsion-free, or $\Gamma(\mathtt{Y},\mathtt{Z})=\Gamma(\mathtt{Z},\mathtt{Y})$ for two vector fields $\mathtt{Y},\mathtt{Z}$, and metric compatible ($\mathrm{D}_{\mathtt{Y}}\langle\mathtt{Z},\mathtt{Z}\rangle_{\mathbf{g}}=2\langle\mathtt{Z},\nabla_{\mathtt{Y}}\mathtt{Z}\rangle_{\mathbf{g}}$.)
A geodesic from $X$ with initial velocity $\xi\in T_{X}\mathcal{M}$ is a curve $\gamma$ on the manifold satisfying the equation $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$, with the initial value condition $\gamma(0)=X,\dot{\gamma}(0)=\xi$. The map $v\mapsto\gamma(1)$ is the Riemannian exponential map, which is a different notion than Lie group exponential [9, section 2.1, chapter 4]. A central theme that we will exploit is for some special metrics, the two concepts of exponential are related.
If $\eta\in T_{X}\mathcal{M}$ is another tangent vector at $X$, then the parallel transport of $\eta$ along the geodesic $\gamma(t)$ from 0 to $t$ is the vector $\mathtt{T}_{0}^{t}(\gamma)\eta=\Delta(t)$, where $\Delta$ is a vector field along the curve $\gamma$, that means for each $0\leq s\leq t$, $\Delta(s)$ is tangent to $\mathcal{M}$ at $\gamma(s)$. Additionally, $\Delta$ must satisfy the transport equation
@@
In the next section, we will define matrix groups, a major topic of this paper. It also provides background for quotient manifolds as examples of submersion.
@@ 2.3. Matrix Lie groups, Lie algebra, left-invariant metrics, and matrix quotient spaces
A Lie group is a differentiable manifold with a differentiable group structure. We are interested in subgroups of the general linear group $\mathcal{E}=\operatorname{GL}(n)$, which is the (multiplication) group of invertible matrices in $\mathbb{R}^{n\times n}$. Thus, for such a group $\mathcal{G}$, $\operatorname{I}_{n}\in\mathcal{G}$ and $G_{1}^{-1}G_{2}\in\mathcal{G}$ if $G_{1},G_{2}\in\mathcal{G}$.
We also consider cartesian products of matrix Lie groups, equipping the product $\mathcal{G}\times\mathcal{H}$ for two matrix Lie groups $\mathcal{G}\in\mathbb{R}^{n_{1}\times n_{1}},\mathcal{H}\in\mathbb{R}^{n_{2}\times n_{2}}$ with the component-wise multiplication $(G_{1},H_{1})(G_{2},H_{2})=(G_{1}G_{2},H_{1}H_{2})$ and inversion. This is equivalent to embedding $\mathcal{G}\times\mathcal{H}$ in a block diagonal matrix subgroup of $\mathbb{R}^{(n_{1}+n_{2})\times(n_{1}+n_{2})}$, where $\mathcal{G}$ is embedded in the top-left $n_{1}\times n_{1}$-block and $\mathcal{H}$ with the bottom-right $n_{2}\times n_{2}$ block. This realization extends to tangent spaces and Lie algebras discussed below.
@@
The tangent space $\mathfrak{g}$ of $\mathcal{G}$ at $\operatorname{I}=\operatorname{I}_{n}$ is a Lie subalgebra of $\mathbb{R}^{n\times n}$. Recall a vector subspace $\mathfrak{g}$ of the square matrix space $\mathcal{E}=\mathbb{R}^{n\times n}$ is called a Lie subalgebra if it is closed under the Lie bracket, that means for two matrices $\mathtt{A},\mathtt{B}$ in $\mathfrak{g}$,
(5) $\mathtt{A},\mathtt{B}]:=\mathtt{A}\mathtt{B}-\mathtt{B}\mathtt{A}\in\mathfrak{g}.$
@@
The Lie bracket is anticommutative and satisfies the Jacobis identity [9, Chapter 3]. The matrix exponential map $\operatorname{expm}$ maps $\mathfrak{g}$ to $\mathcal{G}$. Conversely, given a Lie subalgebra $\mathfrak{g}$ of $\mathbb{R}^{n\times n}$, we can construct a connected Lie group with Lie algebra $\mathfrak{g}$. In this instant, $\operatorname{expm}$ corresponds to the Lie exponential map of a subgroup of $\operatorname{GL}(n)$. We have
(6) $\operatorname{Ad}_{\operatorname{expm}(\mathtt{Z})}\mathtt{X}=\operatorname{expm}(\mathtt{Z})\mathtt{X}\operatorname{expm}(-\mathtt{Z})=\operatorname{expv}(\operatorname{ad}_{\mathtt{Z}},\mathtt{X}),$
@@
where $\operatorname{ad}_{\mathtt{Z}}:\mathtt{X}\mapsto[\mathtt{Z},\mathtt{X}]$ and $\operatorname{expv}$ is the exponential action in appendix A [9, Chapter 3]. From this, $\operatorname{expm}(\mathtt{Z})\mathtt{X}\operatorname{expm}(-\mathtt{Z})$ is in $\mathfrak{g}$ for $\mathtt{Z},\mathtt{X}\in\mathfrak{g}$. The tangent space of $\mathcal{G}$ at $X\in\mathcal{G}$ could be identified with $X\mathfrak{g}$, or $\xi\in T_{X}\mathcal{G}$ if and only if $X^{-1}\xi\in\mathfrak{g}$. Thus, if $\langle,\rangle$ is a nondegenerate form on $\mathfrak{g}$, we can define a pseudo-Riemannian metric on $\mathcal{G}$ by defining for two tangent vectors $\xi,\eta$ at $X$ $\langle\xi,\eta\rangle_{X}=\langle X^{-1}\xi,X^{-1}\eta\rangle$, which gives $\mathcal{G}$ a left-invariant pseudo-Riemannian metric.
For a left-invariant metric, left multiplication is an isometry, a smooth, invertible map between two manifolds preserving the metric (between the same manifold in this case). A metric is right-invariant if for all $X_{1}\in\mathcal{G}$, the map $R_{X_{1}}:X\mapsto XX_{1},X\in\mathcal{G}$ is an isometry, and it is bi-invariant if it is both left and right invariant, and this is equivalent to $\operatorname{Ad}_{\mathcal{G}}$-invariant on $\mathfrak{g}$, $\langle K\mathtt{X}K^{-1},K\mathtt{X}K^{-1}\rangle=\langle\mathtt{X},\mathtt{X}\rangle$ for $X\in\mathfrak{g},K\in\mathcal{G}$. From [33, Proposition 11.9], if $\mathcal{G}$ is connected, bi-invariance is equivalent to the cyclic invariant condition (note the items in the brackets below preserve the cyclic order)
(7) $\langle\mathtt{A},[\mathtt{B},\mathtt{C}]\rangle=\langle\mathtt{B},[\mathtt{C},\mathtt{A}]\rangle=\langle\mathtt{C},[\mathtt{A},\mathtt{B}]\rangle,\mathtt{A},\mathtt{B},\mathtt{C}\in\mathfrak{g}.$
From the same proposition, bi-invariant pseudo-Riemannian metrics on Lie groups have the important property that geodesics are given by the Lie exponential map.
The concepts of homogeneous spaces, coset manifolds, group actions (both right and left), and isotropy groups are covered in [33, Chapters 9 and 11].
We call a pseudo-Riemannian space $\mathcal{M}$ a reductive space [33, chapter 11] if it is isometric to a quotient $\mathcal{G}/\mathcal{K}$, for Lie groups $\mathcal{G}$ and $\mathcal{K}$ with Lie algebras $\mathfrak{g}$ and $\mathfrak{k}$ such that there is a subspace $\mathfrak{m}\subset\mathfrak{g}$ complementary to $\mathfrak{k}\subset\mathfrak{g}$ with $\mathfrak{m}$ is $\operatorname{Ad}_{\mathcal{K}}$-invariant ($K\mathfrak{m}K^{-1}\subset\mathfrak{m}$ for $K\in\mathcal{K}$, and when $\mathcal{K}$ is connected, this is equivalent to $\mathfrak{k},\mathfrak{m}]\subset\mathfrak{m}$). From [20, Proposition 10.3.1], there is a one-to-one correspondence between $\mathcal{G}$-invariant pseudo-Riemannian metric on $\mathcal{G}/\mathcal{K}$ and $\operatorname{ad}(\mathfrak{k})$-invariant non-degenerate symmetric bilinear form $\langle,\rangle$ on $\mathfrak{m}$. That means
(8) $\langle[\mathtt{K},\mathtt{M}_{1}],\mathtt{M}_{2}\rangle=\langle\mathtt{M}_{1},[\mathtt{K},\mathtt{M}_{2}]\rangle\text{ for }\mathtt{K}\in\mathfrak{k},\mathtt{M}_{1},\mathtt{M}_{2}\in\mathfrak{m}.$
Further, it is naturally reductive if the form satisfies the cyclic condition
(9) $\langle\mathtt{M}_{1},\mathcal{P}_{\mathfrak{m}}[\mathtt{M}_{2},\mathtt{M}_{3}]\rangle=\langle\mathtt{M}_{2},\mathcal{P}_{\mathfrak{m}}[\mathtt{M}_{3},\mathtt{M}_{1}]\rangle\text{ for }\mathtt{M}_{1},\mathtt{M}_{2},\mathtt{M}_{3}\in\mathfrak{m}.$
Here, $\mathcal{P}_{\mathfrak{m}}$ is the projection to $\mathfrak{m}$ (using the fact that $\mathfrak{k}$ and $\mathfrak{m}$ are complementary), the term $\mathcal{P}_{\mathfrak{m}}(\mathtt{M}_{1})$ is typically written $(\mathtt{M}_{1})_{\mathfrak{m}}$ in the differential geometric literature, and $\mathcal{P}_{\mathfrak{m}}([\mathtt{M}_{3},\mathtt{M}_{1}])$ is typically written $\mathtt{M}_{1},\mathtt{M}_{3}]_{\mathfrak{m}}$. Here, we follow the matrix convention.
In the above, the form $\langle,\rangle$ needs to be defined on $\mathfrak{m}$ only. We do not assume a metric on $\mathfrak{k}$, or we have much freedom to extend $\langle,\rangle$ to a left-invariant metric on $\mathcal{G}$, see [33, Lemma 11.24]. However, a construction that is often used is if $\mathfrak{g}$ itself is equipped with a nondegenerate bi-invariant form $\langle,\rangle$ such that its restriction to $\mathfrak{k}$ is nondegenerate, we can take $\mathfrak{m}$ to be the orthogonal complement to $\mathfrak{k}$ in lemma 1. Then, the restriction to $\mathfrak{m}$ is also nondegenerate, and both eq. 8 and eq. 9 are satisfied, giving us a naturally reductive metric on $\mathcal{M}=\mathcal{G}/\mathcal{K}$. This is essentially the construction that we use for the nonsingular case in the proof of theorem 1.
As mentioned in [33, chapter 11.22], naturally reductive metrics could be considered a generalization of the bi-invariant metrics to a quotient subspace. It has a simple expression for the Levi-Civita connection [20, Theorem 3.3] and geodesics. In [39, Proposition 2.12], the author showed if $\mathcal{G}$ is given a bi-invariant metric (which the proof assumes), the parallel transports on $\mathcal{G}/\mathcal{K}$ are also given by a simple equation. This will be the result we use, although we originally derived our results by working directly with the transport equation.
3. The Trace form as a bi-invariant symmetric bilinear form and transposable subalgebras
@@
For matrix Lie algebras, the most important examples of bi-invariant symmetric bilinear forms are trace forms [33, Lemma 11.6]. The trace form is the bilinear map on $\mathbb{R}^{n\times n}$, mapping $\mathtt{A},\mathtt{B}\in\mathbb{R}^{n\times n}$ to the scalar $\langle\mathtt{A},\mathtt{B}\rangle_{\mathscr{K}}:=\operatorname{Tr}\mathtt{A}\mathtt{B}$ (the subscript $\mathcal{K}$ is in honor of Killing, who defined the closely related Killing form). Using the commutative property of trace, we have
(10) $\begin{gathered}\operatorname{Tr}(K\mathtt{B}K^{-1})\mathtt{C}=\operatorname{Tr}\mathtt{B}(K^{-1}\mathtt{C}K)\quad\text{ for }K\in\operatorname{GL}(n),\mathtt{B},\mathtt{C}\in\mathbb{R}^{n\times n},\\ \operatorname{Tr}[\mathtt{A},\mathtt{B}]\mathtt{C}=\operatorname{Tr}\mathtt{B}[\mathtt{C},\mathtt{A}]\quad\text{ for }\mathtt{A},\mathtt{B},\mathtt{C}\in\mathbb{R}^{n\times n},\end{gathered}$
@@
or the trace form is bi-invariant on $\mathbb{R}^{n\times n}$ and on all matrix Lie subalgebras. As any strictly upper-triangular matrix $\mathtt{A}$ has $\operatorname{Tr}\mathtt{A}^{2}=0$, the trace form is not positive definite. A symmetric matrix $\mathtt{A}$ has $\operatorname{Tr}\mathtt{A}^{2}\geq 0$, while an antisymmetric matrix has $\operatorname{Tr}\mathtt{A}^{2}\leq 0$. However, if it is nondegenerate on a Lie algebra $\mathfrak{g}$ of a Lie group $\mathcal{G}$, the trace form induces a bi-invariant pseudo-Riemannian metric on $\mathcal{G}$ with closed-form geodesics.
@@
The Frobenius form, or Frobenius inner product $(\mathtt{A},\mathtt{B})\mapsto\langle\mathtt{A},\mathtt{B}\rangle_{\mathscr{F}}:=\operatorname{Tr}\mathtt{A}\mathtt{B}^{\mathsf{T}}$ of two matrices is positive-definite. However, in general, it is not bi-invariant and may not induce a metric with closed-form geodesics. For subspaces that transpose to themselves, the positive-definite property of Frobenius forms can be used to prove nondegeneracy of subspaces equipped with the bi-invariant trace form.
@@
In the following, transposability implies nondegeneracy. This may be well-known in the reductive Lie algebra literature, but we do not have an exact reference.
@@
If $\mathfrak{g}$ is transposable then the trace form is nondegenerate on $\mathfrak{g}$. Let $\mathfrak{v}\subset\mathfrak{g}$ be a transposable subspace. Let $\mathfrak{v}_{{}_{\perp}}$ be the orthogonal complement of $\mathfrak{v}$ under the trace form $\langle\mathtt{A},\mathtt{B}\rangle_{\mathscr{K}}=\operatorname{Tr}\mathtt{A}\mathtt{B}$
(12) $\mathfrak{v}_{{}_{\perp}}=\{\mathtt{A}\in\mathfrak{g}|\;\operatorname{Tr}\mathtt{A}\mathfrak{v}=\{0\}\}.$
Then $\mathfrak{v}_{{}_{\perp}}$ is also transposable and is the orthogonal complement under the Frobenius form $\langle\mathtt{A},\mathtt{B}\rangle_{\mathscr{F}}=\operatorname{Tr}\mathtt{A}^{\mathsf{T}}\mathtt{B}$, and vice-versa. The projection $\mathcal{P}_{\mathfrak{v}}$ of $\mathfrak{g}$ onto $\mathfrak{v}$ under $\langle,\rangle_{\mathscr{K}}$ or $\langle,\rangle_{\mathscr{F}}$ is identical and commutes with transpose, $\mathcal{P}_{\mathfrak{v}}(\mathtt{A}^{\mathsf{T}})=\mathcal{P}_{\mathfrak{v}}(\mathtt{A})^{\mathsf{T}}$ for $\mathtt{A}\in\mathfrak{g}$.